OpenStudy (anonymous):

Evaluate (9^9) (9^-5)

OpenStudy (lgbasallote):

9^9-5 = 9^4

OpenStudy (anonymous):

9^4

OpenStudy (anonymous):

(a^x)*(a^y)=a^(x+y).You can see why.Use that to get answers